OpenStudy (anonymous):

Susie has chosen a set of vases for her wedding registry. The vases are cylindrical and their dimensions (in inches) are as follows. • Vase A: h = 14 and r = 3 • Vase B: h = 7 and r = 6 Which of the following statements about the volumes of the vases is true? The volume of vase B is twice the volume of vaseA. The volume of vase B is four times the volume of vase A. The volume of vase A is twice the volume of vaseB. The volumes of the two vases are equal.

OpenStudy (jordanloveangel):

@iGreen. @sleepyhead314

OpenStudy (anonymous):

Plug them into the formula: \(V = \pi r^2 h\) Vase A: \(V = (3.14)(3^2)(14)\) \(V = (3.14)(9)(14)\) Multiply 3.14 * 9 * 14 Vase B: \(V = (3.14)(6^2)(7)\) \(V = (3.14)(36)(7)\) Multiply 3.14 * 36 * 7

OpenStudy (anonymous):

Tell me what you get for both of them.

OpenStudy (anonymous):

just a sec my sis took my calculator

OpenStudy (anonymous):

Kk

OpenStudy (anonymous):

vase a. 395.64 vase b. 791.28

OpenStudy (anonymous):

Yes, now divide 791.28 / 395.64.

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Yep, so Vase B is twice the volume of Vase A.